The length of the side of a triangle are in a ratio of 3:3:5. The perimeter of the triangle is 44cm. Find the length of the sides.
Answer by DrBeeee(684)

This problem allows me to introduce you to "Nicole's n" (named after one of my brilliant students). It really is the application of equivalent fractions. For example 2/3 = 4/6 or 8/9 = 8n/9n. In your problem the proportion is given by
3:3:5. Nicole says that's easy, just use 3n:3n:5n to represent the sides of the triangle. PLEASE remember this very simple technique.
The perimeter of the triangle is the sum of the three sides or
3n + 3n + 5n = perimeter = 44 which simplifies to
11n = 44 or
n = 4
The three sides are 3n, 3n, 5n or
3*4, 3*4, 5*4 or
12, 12, 20
Is this correct?
Is (12+12+20 = 44)?
Is (44 = 44)? Yes
Answer: The three sides of the triangle are 12cm, 12cm, and 20cm.